### v2

@@ Abstract
Artificial intelligence (AI)–driven models have the potential to revolutionize weather forecasting, but still rely on initial conditions generated by costly Numerical Weather Prediction (NWP) systems. Although recent end-to-end forecasting models attempt to bypass NWP systems, these methods lack scalable assimilation of new types of observational data. Here, we introduce XiChen, an observation-scalable fully AI-driven global weather forecasting system, wherein the entire pipeline, from Data Assimilation (DA) to medium-range forecasting, can be accomplished within only 15 seconds. XiChen is built upon a foundation model that is pre-trained for weather forecasting and subsequently fine-tuned to serve as both observation operators and DA models, thereby enabling the scalable assimilation of conventional and raw satellite observations. Furthermore, the integration of Four-Dimensional Variational (4DVar) knowledge ensures XiChen to achieve DA and medium-range forecasting accuracy comparable to operational NWP systems, with skillful forecasting lead time beyond 8.75 days. A key feature of XiChen is its ability to maintain physical balance constraints during DA, enabling observed variables to correct unobserved ones effectively. In single-point perturbation DA experiments, XiChen exhibits flow-dependent characteristics similar to those of traditional 4DVar systems. These results demonstrate that XiChen holds strong potential for fully AI-driven weather forecasting independent of NWP systems.
@@
fully AI-driven global weather forecasting, deep learning, four-dimensional variational, data assimilation, foundation model
@@ 1 Introduction
A century ago, Abbe [abbe1901physical] and Bjerknes [bjerknes1904problem] conceptualized weather forecasting as an initial value problem, where forecasts could be derived by starting from initial conditions estimated from observations and applying physical laws. At that time, atmospheric observations were extremely limited, computers did not exist, and the predictability of weather remained largely unknown. Thus, their proposition was remarkably ambitious [bauer2015quiet]. But today, advancements in our understanding of atmospheric mechanisms, the continuous accumulation of observational data, and significant improvements in computational power have collectively enhanced the accuracy of Numerical Weather Prediction (NWP). NWP consists of two core components. The first is Data Assimilation (DA), the process of integrating observations from satellites, meteorological stations, ships, and other sources to estimate the most likely state of the atmosphere and provide accurate initial conditions for forecasting [bauer2015quiet, gettelman2022future]. The second component is the forecast model, which traditionally simulates the spatial-temporal evolution of atmospheric variables by solving a set of complex Partial Differential Equations (PDEs) [bauer2015quiet, gettelman2022future]. Over the past several decades, NWP has undergone a profound technological transformation [bauer2015quiet], making it an indispensable part of modern society [gettelman2022future].
However, the deceleration of Moore’s Law, the increasing complexity of modern NWP systems, and substantial improvements in resolution have made it difficult to produce timely forecasts within limited computational resources [gettelman2022future, bi2023accurate]. These constraints have consequently hindered further progress in NWP. In recent years, the rapid development of Artificial Intelligence (AI) technology has triggered a transformative revolution in weather forecasting [ling2024improving]. On dedicated computing hardware, such as Graph Processing Units (GPUs), AI-based models—such as FourCastNet [kurth2023fourcastnet], Pangu-Weather [bi2023accurate], GraphCast [lam2023learning], FengWu [ailabchen2025fengwu], Fuxi [fudanchen2023fuxi], and AIFS [lang2024aifs]—achieved acceleration rates three to four orders of magnitude faster than operational NWP systems while maintaining comparable forecasting performance. However, these AI models still rely on the (re)analysis fields produced by NWP systems to generate forecasts and are far from a fully AI-driven global weather forecasting system [chen2023towards, xu2024fuxida, wang2024accurate]. For instance, the analysis fields used by current AI-based forecasting models (e.g., ECMWF’s AIFS [lang2024aifs]) are products generated through the Four-Dimensional Variational (4DVar) method [courtier1994strategy]. The 4DVar method employs optimal control theory to assimilate over 10 million observational data points within each 12-hour Data Assimilation Window (DAW) [lupu2017data]. It effectively integrates indirect satellite observations, addressing gaps in direct atmospheric observations that are challenging to capture using sparse conventional observations [lupu2017data]. Nevertheless, the 4DVar method in operational NWP systems requires hours of computation on supercomputers [bauer2020ecmwf], restricting the volume of observations that can be assimilated within a limited time frame, with only 5-10% of the available satellite data being used [bauer2015quiet]. Additionally, 4DVar requires meticulous tuning and precise formulation of a complex background error covariance matrix [gustafsson2018survey]. It also depends on sophisticated observation operators [liang2023machine], tangent linear models, and adjoint models [gustafsson2018survey], all of which demand extensive development time and expertise from highly trained specialists.
In this era of rapid advancements in AI, a bold proposition has emerged: the development of data-driven systems based on AI technology to forecast future weather states [schultz2021can, mcnally2024data] without dependence upon the initial fields from NWP systems. Notably, models such as Aardvark [allen2025end] and GraphDOP [alexe2024graphdop] have successfully generated forecasts directly from raw observations, demonstrating promising performance. However, the natural sparsity of observations and the lack of rigorous quality control pose significant challenges to achieving forecast accuracy comparable to operational NWP systems. Recent studies have also explored integrating pre-trained AI-based weather forecasting models into traditional 4DVar frameworks by leveraging automatic differentiation to create hybrid 4DVar systems [xiao2024fengwu4dvar, li2024fuxien4dvar, fan2025latent4dvar]. Nevertheless, these efforts often rely on idealized models or simulated observations. Even when real-world conventional observations are employed, these approaches still overlook satellite observations, which are critical for operational NWP systems. On the other hand, some researchers have also ambitiously pursued the development of data-driven DA systems [cintra2016tracking, wu2021fast, arcucci2021deep, huang2024diffda, wang2024four], such as FengWu-Adas [chen2023fengwuadas], FuXi-DA [xu2024fuxida], DiffDA [huang2024diffda], and 4DVarFormer [wang2024accurate, wang2024benchmark]. These models utilize modern neural network architectures to assimilate conventional and satellite observations, integrating them with AI-based weather forecasting models to create fully AI-driven global weather forecasting systems. Despite these advancements, they can only assimilate observation types encountered during training, necessitating retraining and architectural refinement of the entire model to incorporate new observation types. This limitation restricts their scalability compared to traditional DA systems. Consequently, developing fully AI-driven global weather forecasting systems suitable for operational implementation remains an unresolved challenge.
Here, we propose XiChen (Chinese: 羲辰), an observation-scalable fully AI-driven global weather forecasting system. XiChen is constructed by a foundation model [awais2025foundation] trained on diverse weather data to replace core components of the NWP pipeline. The foundation model is initially pre-trained on ERA5 [hersbach2020era5] data to learn fundamental atmospheric dynamical representations, serving as the medium-range weather forecasting model. Subsequently, it is fine-tuned to incorporate the observation operators and the DA capabilities. By incorporating the 4DVar knowledge into the DA model, specifically by leveraging the gradient of the AI-driven 4DVar cost function as model inputs, we propose a cascaded sequential DA framework to effectively assimilate both conventional and raw satellite observations for initializing accurate global weather forecasting. This cascaded DA framework shows the potential to address the compatibility and robustness issues associated with the future integration of new observations by fine-tuning. In our experiments, XiChen assimilates conventional observations from the Global Data Assimilation System (GDAS), raw satellite data from the Microwave Humidity Sensor (MHS) and Advanced Microwave Sounding Unit-A (AMSU-A), as well as wind field retrievals from the Advanced Scatterometer (ASCAT) and satellite-derived winds (SATWND). Although XiChen operates on a global grid with a resolution of 1.40625, its accuracy is comparable to that of operational NWP systems across multiple variables in DA and medium-range weather forecasting tasks. Furthermore, the entire pipeline of XiChen, from DA to 10-day medium-range forecasting, can be accomplished within only 15 seconds using a single A100 GPU, achieving over 400-fold acceleration compared to the operational NWP system [bauer2020ecmwf], with a skillful forecast lead time of over 8.75 days, surpassing the Global Forecasting System (GFS). A distinctive capability of XiChen lies in its ability to maintain physical balance constraints during DA tasks. This enables the system to utilize observed variables to effectively correct other unobserved but interrelated ones, ensuring internal consistency. In single-point perturbation DA experiments, XiChen exhibited flow-dependent characteristics comparable to those observed in the traditional 4DVar system [zhu2022four]. A comprehensive evaluation is presented for three key aspects: 1) the impacts of various DA observation configurations on medium-range forecasting; 2) the effects of 4DVar knowledge on both DA and forecasting; and 3) XiChen’s performance on the Tropical Cyclone (TC) track task. The promising findings highlight that XiChen offers a novel avenue for advancing fully AI-driven global weather forecasting through fine-tuning to scalably integrate new observations in the future.
@@ 2 Results
Our experiments evaluated two key aspects of XiChen’s performance: (1) the accuracy and stability of its one-year DA cycle, and (2) the skill of its 10-day medium-range weather forecasts. The DA process employed a cascade framework, sequentially assimilating GDAS prepbufr data, followed by MHS, ASCAT, AMSU-A, and SATWND observations. For the DA cycle experiment, XiChen was initialized at 00:00 UTC on 1 January 2023. The background field was generated by initializing the XiChen forecast model with ERA5 reanalysis data from 00:00 UTC on 29 December 2022, followed by a 3-step rollout with a 24-hour lead time. This forecast field served as the initial condition for the DA cycle, which ran for one year with assimilations at 00:00 and 12:00 UTC daily. We employ a 12-hour DAW wherein observations at 0, 3, 6, and 9 hours after the analysis time are assimilated to initialize the forecasting model for generating the background field for the subsequent DA cycle. To improve the quality of the analysis field, we constructed an ensemble DA using 11 ensemble members within the 12-hour DAW. The weight of each ensemble member was determined based on its distance from the GDAS prepbufr. For detailed information on the ensemble averaging configuration, refer to Section 4. Furthermore, we implement a 6-hour DAW to establish the initial conditions for medium-range forecasting. This computational process is executed within 3 seconds utilizing a single A100 GPU. This dual DA framework ensures the fidelity of the assimilation process for medium-range predictions. Further details of the cascade process and DA cycle are illustrated in Figures S1(a) and S1(b). The weather forecasting model, DA models, and observation operators share the same underlying model architecture and are trained using different data. Details regarding this architecture are provided in Section 4. For medium-range weather forecasting, we adopted configurations and evaluation metrics from WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark] to generate initial fields and evaluate XiChen’s 10-day forecasts. For details, please see Section 4.8 and Figure S1(c).
Additionally, we evaluated XiChen’s TC assimilation and forecasting capabilities. TC tracks were identified using the TempestExtremes [ullrich2021tempestextremes] tracker, configured as in NeuralGCM [kochkov2024neural], and compared against the International Best Track Archive for Climate Stewardship (IBTrACS) and ERA5. Tracker parameters are detailed in Table S9, and TC track results from forecasts at different lead times are shown in Figures S9. Finally, we present visualizations of the forecasting results produced by XiChen, IFS HRES, and GFS in Figures S10 to S15.
2.1 One year DA Cycle
The DA performance of XiChen was systematically evaluated against State-Of-The-Art (SOTA) operational NWP systems, specifically the IFS HRES and GFS. These variables were carefully selected to provide a comprehensive assessment of the system’s predictive capabilities across multiple atmospheric levels and meteorological parameters.
Figure 1 presents the time series of globally averaged annual Root Mean Square Error (RMSE) for the DA cycle results of XiChen, using the ERA5 data as the reference. The RMSE values are compared against those of the operational analysis fields from the GFS and IFS HRES models. Each subplot corresponds to a specific variable, with skill (y-axis) plotted at 12-hour intervals over a 365-day forecast horizon (x-axis). The original data is displayed with reduced opacity for clarity, while solid lines represent values smoothed using an Exponential Moving Average (EMA) with a 15-point window. XiChen demonstrates stable performance throughout the year‑long DA cycle and effectively leverages limited observations to estimate key atmospheric variables. For surface variables with relatively dense direct measurements, such as T2M, U10, and V10, XiChen achieves lower RMSE than GFS. XiChen also attains RMSE values comparable to or lower than those of GFS for Z300, Z500, Z700, and T850. In contrast, XiChen exhibits larger errors for T300, T500, U300, U500, U700, Q300, Q500, and Q700, largely due to the scarcity of observations at upper levels. Occasional increases in RMSE, as seen in Figure 1, are primarily attributed to missing observations. However, XiChen effectively addresses these issues through subsequent assimilation processes.
Figure 1: Comparison of the average latitude RMSE of the one-year DA cycle using XiChen (red lines) as well as the operational analysis fields of GFS (yellow lines) and IFS HRES (blue lines) using testing data during 2023. The analysis encompassed 16 distinct variables, including geopotential heights at 300 hPa (Z300), 500 hPa (Z500), and 700 hPa (Z700); 2-meter temperature (T2M); temperatures at 300 hPa (T300), 500 hPa (T500), 700 hPa (T700), and 850 hPa (T850); u- and v-components of the wind at 10 meters (U10 and V10, respectively); u-components of the wind at 300 hPa (U300), 500 hPa (U500), and 700 hPa (U700); and specific humidity at 300 hPa (Q300), 500 hPa (Q500), and 700 hPa (Q700). All analysis fields are evaluated against the ERA5 reanalysis dataset.
2.2 Medium-range weather forecasting
@@
The primary criterion for evaluating the fully AI-driven global weather forecasting system lies in its efficacy in generating medium-range global weather forecasts. In this investigation, we evaluate the medium-range global weather forecasts produced at 6-hour intervals by XiChen. The forecast output is systematically compared against reanalysis data from the ERA5 dataset. The temporal initialization of these analysis fields aligns with the configurations established by WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark], ensuring consistency with prior benchmarks. XiChen’s forecasting performance is evaluated against SOTA operational NWP systems, specifically the IFS HRES and GFS models.
@@
Figure 2 illustrates the averaged latitude-weighted RMSE and Anomaly Correlation Coefficient (ACC), both as a function of forecast lead times over 10 days. Each subplot corresponds to a variable and skill (y-axis) is plotted at 6-hour steps over 10-day horizons (x-axis). Rows 1 and 3 show RMSEs, rows 2 and 4 show the ACCs. When using data from the ERA5 dataset as the initial fields, XiChen outperforms IFS HRES and GFS in all evaluated variables (the red dashed lines). XiChen achieves a skillful weather forecasting lead time exceeding 10.25 days (with ACC of Z500 >0.6>0.6). This performance rivals SOTA medium-range global weather forecasting models trained on 0.25-degree high-resolution datasets—Pangu [bi2023accurate], GraphCast [lam2023learning], FengWu [ailabchen2025fengwu], and FuXi [fudanchen2023fuxi].
@@
Figure 2: Comparison of the average latitude RMSE (first and third rows) and ACC (second and fourth rows) of the 10-day medium-range weather forecasting using XiChen (red lines) as well as the operational forecasting results of GFS (yellow lines) and IFS HRES (blue lines) using testing data during 2023. The evaluation encompasses eight meteorological variables: geopotential height at 500 hPa (Z500) and 850 hPa (Z850), 2-meter temperature (T2M), temperature at 850 hPa (T850), Mean Sea Level Pressure (MSLP), 10-meter zonal (10U) and meridional (10V) wind components, as well as specific humidity at 500 hPa (Q500). All analysis fields are evaluated against the ERA5 reanalysis dataset.
@@
When using XiChen’s self-generated initial conditions, XiChen achieved a skillful weather forecasting lead time exceeding 8.75 days, surpassing the GFS system (8.25 days). Notably, for key variables such as T2M, T850, MSLP, U10, and V10, the fully AI-driven global weather forecasts of XiChen deliver lower RMSEs compared to the GFS for nearly the entire 10-day lead time. The RMSEs for all variables approach or even surpass the performance of the IFS HRES as the lead time increases. For Q500, XiChen exceeds the GFS after 2 days and the IFS HRES after 4 days. Although the forecast error for Z500 remains higher than that of both the GFS and the IFS HRES during the first 6 days, XiChen surpasses the GFS after 7 days and outperforms IFS HRES after 8 days. This achievement is particularly noteworthy considering that XiChen was trained using only 12 years of ERA5 data at a 1.40625 resolution. Moreover, XiChen assimilated a limited subset of observations used in operational NWP systems–including GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND observations–significantly fewer than those used by GFS and IFS HRES. XiChen’s RMSE surpasses that of IFS HRES at larger lead times, which may be because its forecasts tend to favor smoothing, as evidenced by the energy spectra presented in Figure S5. This tendency is a common characteristic observed in contemporary AI-based weather forecasting models [lam2023learning, ailabchen2025fengwu, fudanchen2023fuxi]. For comprehensive insights into the 10-day medium-range forecast RMSE scorecards across all variables for Xichen versus GFS, and XiChen versus IFS HRES, kindly consult the Supplementary Text, as well as Figures S6 and S7.
2.3 Contribution of assimilating different observations

### v3

@@ Abstract
Machine Learning (ML) has shown great promise in revolutionizing weather forecasting, yet most ML systems still rely on initial conditions generated by Numerical Weather Prediction (NWP) systems. End-to-end ML models aim to eliminate this dependency, but they often rely on observation-specific encoders and require redesign or retraining when observation sources change, thereby limiting their operational robustness. Here, we introduce XiChen, a global weather observation-to-forecast ML system via four-dimensional variational (4DVar) gradient-guided flexible assimilation. We demonstrate that the gradient of the 4DVar cost function serves as a physically grounded interface that maps heterogeneous observations into a common state space. This novel formulation enables XiChen to flexibly assimilate diverse conventional and raw satellite observations while preserving physical consistency. Experiments show that the system achieves forecasting metrics competitive with operational NWP systems. This work provides a practical and physically consistent route toward operational ML-based global weather forecasting systems with heterogeneous and evolving observations.
@@
data-to-forecast ML system, four-dimensional variational, data assimilation, foundation model, heterogeneous observation
@@ 1 Introduction
Accurate medium-range global weather forecasting is a cornerstone for transportation, disaster early warning systems, and renewable energy integration. Over the past decades, physics-based Numerical Weather Prediction (NWP) systems have achieved remarkable progress [bauer2015quiet], driven by synergistic advances in computational power [bi2023accurate], Data Assimilation (DA) techniques [wang2000data, bannister2017review], observational infrastructures [tatem2008fifty], and physical understanding [gettelman2022future]. However, as we enter an era of explosive growth in global satellite and sensor data streams, traditional NWP systems are approaching practical computational limits [bauer2021digital, ecmwf10years]. This limitation suggests that incremental improvements to existing NWP pipelines are insufficient to keep pace with the influx of data from future observational systems. Consequently, there is a compelling motivation to transition toward fundamentally new paradigms that can fully exploit data-rich environments.
Recent advances in Machine Learning (ML) have demonstrated substantial potential for simulating complex weather systems without the costly step-by-step numerical integration of partial differential equations [bi2023accurate, lam2023learning, fudanchen2023fuxi, ailabchen2025fengwu]. However, most ML-based weather forecasting models lack inherent DA capabilities and typically rely on initial conditions provided by traditional NWP systems [chen2023fengwuadas, xiao2024fengwu4dvar, wang2024accurate], preventing the realization of a fully self-contained pipeline from observations to forecasting results. Emerging end-to-end ML models [chen2023fengwuadas, xu2024fuxida, mcnally2024data, alexe2024graphdop, allen2025end, sun2025data] seek to overcome this dependency, but they face a major limitation in adapting to the explosive growth and extreme heterogeneity of global satellite and sensor data [gettelman2022future, bauer2021digital]. Current approaches rely on observation-type-specific encoders tightly coupled to sensor characteristics, requiring redesign or retraining the entire pipeline whenever observation sources are missing, evolved, or are newly introduced [xu2024fuxida, allen2025end, sun2025data]. This limitation is particularly problematic given the rapid evolution of Earth observation systems, where new satellite platforms and sensing modalities are continuously deployed [gettelman2022future]. Consequently, the central challenge in achieving a fully self-contained ML pipeline is the absence of a physically consistent and flexible ML-based DA framework that scales with evolving global observations.
Integrating the four-dimensional variational (4DVar) [carrassi2018data] DA method with ML techniques offers a promising approach to addressing the above challenge. This synergy is driven by the fact that the gradient of the 4DVar cost function encodes the sensitivity of atmospheric states to observational discrepancies, reflecting the physical constraints imposed by observations on the initial state through the forecasting model. Recent ML-based 4DVar frameworks leverage the automatic differentiation capabilities of ML-based weather forecasting models to replace the tangent-linear and adjoint models traditionally required in operational 4DVar systems [xiao2024fengwu4dvar, li2024fuxien4dvar, wang2024accurate]. However, existing research has primarily utilized simulated conventional observations and has yet to demonstrate scalability to real-world, high-dimensional, and heterogeneous observations. To bridge this gap, it is essential to internalize the 4DVar constraint within ML architectures. This integration can facilitate a unified and physically consistent framework capable of learning complex observation-to-state mappings directly from diverse conventional and raw satellite observations.
Here, we propose XiChen, a flexible ML-based global DA and medium-range global weather forecasting system. To address the challenges posed by data heterogeneity, XiChen leverages a foundation model pre-trained for global weather forecasting, which is subsequently fine-tuned to function as observation operators and DA models. By utilizing the 4DVar gradient as a physically consistent unifying information interface, XiChen transforms multi-source observations into a common state space that the DA model can seamlessly ingest. Within this framework, the 4DVar cost function is formulated using the ML-based weather forecasting model and observation operators, with its gradient relative to the background field computed via automatic differentiation. This approach enables XiChen to facilitate stable, long-term DA cycles by leveraging multi-source heterogeneous real-world observations. Experimental results indicate that the DA process of XiChen achieves physical consistency and flow-dependent behavior, reflecting the core advantages of integrating 4DVar within the ML model. Furthermore, the medium-range global weather forecasting metrics of XiChen are comparable to those of advanced operational NWP systems for key atmospheric variables. Finally, we demonstrate its potential for the continuous, scalable assimilation of heterogeneous observations and the iterative improvement of forecasting metrics.
@@ 2 Results
As illustrated in Figure 1(a), existing data-driven DA methods encounter two primary challenges in handling the complex heterogeneous observation system: (1) end-to-end models fuse observations and background fields in a purely black-box manner, making it difficult to ensure physical consistency; and (2) each observation type typically requires a specific encoder, necessitating a full pipeline retraining to accommodate missing, evolving, or newly introduced observations. To address these limitations, we show that the gradient of the 4DVar cost function serves as a unified information interface for heterogeneous observations. This principle enables XiChen, a physically consistent ML-based system for medium-range global weather forecasting and observation-flexible DA. As illustrated in Figure 1(b) and (c), we implement XiChen by pretraining a foundation model as a medium-range global weather forecasting model and subsequently fine-tuning it as the observation operators and DA models. Specifically, we construct a 4DVar cost function by integrating the forecasting model with the observation operators. The gradient of the 4DVar cost function is then calculated via automatic differentiation to propagate information from heterogeneous observations across the spatial and multi-variable dimensions of the atmospheric state (see Section 4 for details). For details of the DA cycle and medium-range forecasting experiments, see the Supplementary Text, Figure S1(a) and S1(b).
XiChen is evaluated over one year in 2023. The evaluation focuses on four key aspects: (1) single-point perturbation assimilation of satellite radiances to assess the physical consistency of the DA process of XiChen; (2) the skill of XiChen’s 10-day medium-range global weather forecasts; (3) the impact of assimilating different observations on medium-range global weather forecasts; and (4) the influence of satellite assimilation on Tropical Cyclone (TC) forecasting skill. The DA process employs a cascade framework, sequentially assimilating conventional observations from the Global Data Assimilation System (GDAS) prepbufr data, followed by Microwave Humidity Sensor (MHS) brightness, wind field retrievals from the Advanced Scatterometer (ASCAT), Advanced Microwave Sounding Unit-A (AMSU-A) brightness, and satellite-derived winds (SATWND) observations. The weather forecasting model, observation operators, and DA models share the same underlying model architecture and are trained using different data. For medium-range global weather forecasting, we adopted configurations and evaluation metrics from WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark] to generate initial fields and evaluate XiChen’s 10-day forecasts (see Section 4.7 and Figure S1(c) for details).
Figure 1: Overview of our XiChen. (a) A schematic illustration demonstrating the utility of the 4DVar cost function gradient as a unified information interface of observations. (b) The overall framework of “XiChen”: a global weather observation-to-forecast machine ML system via 4DVar gradient-guided flexible assimilation. XiChen integrates background field data, conventional observations, and indirect satellite observations within the Data Assimilation Window (DAW) to perform the DA task. The process begins by computing the 4DVar cost function, and it subsequently determines the gradient with respect to the background field. By combining this gradient with the background field, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. The green box denotes the weather forecasting model, the purple represents the observation operator, and the blue signifies the DA model. XiChen undergoes three stages of training: (c) It is pre-trained as a medium-range global weather forecasting model, with lead time ($\Delta t$) as the conditional input; (d) It is fine-tuned as an observation operator, using meta data of the satellite as the conditional input; (e) It is fine-tuned as a DA model, with the gradient of the 4DVar cost function as the input and the background field as the conditional input.
2.1 Physical consistency of the assimilation
To evaluate whether the XiChen assimilation process exhibits fundamental physical consistency, we designed a single-point perturbation experiment to assess the system’s response to localized observation changes. The evaluation consists of two separate DA runs. The first run used ERA5 as the background field to assimilate raw satellite observations. In the second run, a perturbation is introduced to the raw satellite inputs across an individual observation channel. The differences between the two DA results illustrate the impact of the perturbation on the analysis fields, revealing how observational information propagates through the variables and spatial dimensions. The DA time was set to 00 UTC on July 24, 2023, and perturbation experiments were conducted separately for AMSU-A and MHS. The perturbation location, selected at 20°N, 120°E, was near the path of Typhoon Doksuri during this period. The perturbation magnitude was set to +5 K, and independent experiments were performed for each satellite channel.
Figures 2(a) and 2(b) present the horizontal and vertical distributions of analysis field variations resulting from individually perturbing three AMSU-A channels (5, 6, and 7) and three MHS channels (3, 4, and 5), respectively. For AMSU-A, increases in brightness temperature correspond to rising atmospheric temperatures, consistent with radiative transfer theory, demonstrating that XiChen has effectively learned AMSU-A’s temperature sensitivity. For MHS, increases in brightness temperature correspond to decreases in specific humidity, indicating that XiChen has similarly captured MHS’s sensitivity to atmospheric moisture. In both experiments, the analysis increments exhibit propagating variations horizontally and vertically and display localized, anisotropic structures around the observations that align with the background geopotential field, reflecting the flow-dependent nature of the XiChen system. Additionally, for AMSU-A, anticyclonic wind responses surrounding temperature increments imply that XiChen satisfies certain equilibrium constraints. Overall, these results demonstrate that XiChen adheres to fundamental physical consistency, which is consistent with the traditional 4DVar system [zhu2022four].
Figure 2: Analysis field changes from a 5 K perturbation to AMSU-A and MHS observations. The perturbation was applied at 03 UTC near Typhoon Doksuri (20N, 120E; yellow dot). (a) AMSU-A results (channels 5-7). Top row: horizontal temperature changes (shading, K) and wind differences (vectors, m s-1) at 700, 200, and 200 hPa. Solid contours show background geopotential (m2 s-2). Bottom row: vertical cross-section. (b) MHS results (channels 3-5). Top row: horizontal specific humidity changes (shading, g kg-1) and wind differences (vectors, m s-1) at 300, 500, and 500 hPa. Bottom row: vertical cross-section.
2.2 Medium-range global weather forecasting
@@
As forecasting accuracy is the primary criterion for the ML-based global DA and medium-range global weather forecasting system, we evaluate the performance of XiChen against advanced operational NWP systems. This evaluation is based on global weather forecasts at 6-hour intervals. The temporal initialization of these analysis fields aligns with the configurations established by WeatherBench [rasp2020weatherbench] and DABench [wang2024benchmark].
@@
Figure 3 illustrates the averaged latitude-weighted Root Mean Square Error (RMSE) and Anomaly Correlation Coefficient (ACC), both as a function of forecast lead times over 10 days. XiChen achieves a skillful weather forecasting lead time exceeding 10.25 days (with ACC of Z500 >0.6>0.6). This performance rivals advanced medium-range global weather forecasting models trained on 0.25 high-resolution datasets—Pangu [bi2023accurate], GraphCast [lam2023learning], FengWu [ailabchen2025fengwu], and FuXi [fudanchen2023fuxi].
@@
Figure 3: Comparison of the average latitude RMSE (first and third rows) and ACC (second and fourth rows) of the 10-day medium-range global weather forecasting using XiChen (red lines) as well as the operational forecasting results of GFS (yellow lines) and IFS HRES (blue lines) using testing data during 2023. The evaluation encompasses eight meteorological variables: geopotential height at 500 hPa (Z500) and 850 hPa (Z850), 2-meter temperature (T2M), temperature at 850 hPa (T850), Mean Sea Level Pressure (MSLP), 10-meter zonal (10U) and meridional (10V) wind components, as well as specific humidity at 500 hPa (Q500). All analysis fields are evaluated against the ERA5 reanalysis dataset.
@@
When using XiChen’s self-generated initial conditions, XiChen achieves a skillful weather forecasting lead time exceeding 8.75 days, surpassing the GFS system (8.25 days). Notably, for key variables such as T2M, T850, MSLP, U10, and V10, the global weather forecasts of XiChen deliver lower RMSEs compared to the GFS for nearly the entire 10-day lead time. The RMSEs for all variables approach or even surpass the performance of the IFS HRES as the lead time increases. For RMSE of Q500, XiChen exceeds the GFS after 2 days and the IFS HRES after 4 days. Although the forecast error for Z500 remains higher than that of both the GFS and the IFS HRES during the first 6 days, XiChen surpasses the GFS after 7 days and outperforms IFS HRES after 8 days. This achievement is particularly noteworthy considering that XiChen was trained using only 12 years of ERA5 data at a 1.40625 resolution. Moreover, XiChen assimilated a limited subset of observations used in operational NWP systems, including GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND, significantly fewer than those used by GFS and IFS HRES. XiChen’s RMSE surpasses that of IFS HRES at larger lead times, which may be because its forecasts tend to favor smoothing, as evidenced by the power spectra presented in Figure S6. This tendency is a common characteristic observed in contemporary ML-based weather forecasting models [lam2023learning, ailabchen2025fengwu, fudanchen2023fuxi]. For comprehensive insights into the 10-day medium-range forecast RMSE scorecards across all variables for XiChen versus GFS, and XiChen versus IFS HRES, kindly consult the Supplementary Text, as well as Figures S7 and S8. Furthermore, we present visualizations of the forecasting results produced by XiChen, IFS HRES, and GFS in Figures S12 to S17.
2.3 The impact of assimilating different observations on medium-range global weather forecasts